OpenStudy (anonymous):

Integrate: sin(3t)^4 dt. Definite intregral from 0 to pi.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi} \sin^{4}(3t) dt\]

OpenStudy (dumbcow):

you can use integration by parts, but its a lot of steps there are tables that show integral sin^n(x) Also wolfram provides steps as well http://www.wolframalpha.com/input/?i=integrate+sin^4%283x%29dx

OpenStudy (anonymous):

\[sin^2t = \frac{1 - cos2t}{2}\] Use this to reduce the calculation and algebra and stuff...

OpenStudy (anonymous):

\[\int_0^{\pi} \frac{1}{4}* (1 - cos6t )^2 dt\]

OpenStudy (anonymous):

\[\frac{1}{4}\int _ 0^{\pi} (1 + cos^26t - 2cos6t)dt\]

OpenStudy (anonymous):

Is there a way to do with without the reduction formula and just using trig identities and u-substitution?

OpenStudy (anonymous):

\[cos^2t = \frac{1 + cos2t}{2}\] \[\frac{1}{4} \int_0^{\pi} (1+ \frac{(1 + cos12t)}{4} -2cos6t)dt\] \[\frac{1}{4} \int 1*dt -2cos6t*dt + \frac{1}{16}\int_0^{\pi} 1*dt + cos12t *dt \]

OpenStudy (dumbcow):

yeah, do what ishaan is showing you :)

OpenStudy (anonymous):

\[\left| \frac{t}{4} -\frac{1}{12}sin6t +\frac{t}{16} + \frac{1}{16*12}sin12t\right|_0^{\pi}\]

OpenStudy (anonymous):

Now Algebra ....you can do it ....I may messed up something so do check again the whole integration .... Good Luck

OpenStudy (anonymous):

Ugh my browser was wigging and I couldn't ask questions. The 2nd time you used the half angle formula, why is there a 4 in the denominator?

OpenStudy (anonymous):

ah my bad I told you I may have messed ...I messed it up sorry 2 should be there

OpenStudy (anonymous):

That will make 16 into 8 as I multiplied an extra 2

OpenStudy (anonymous):

So I have 3/8t - 1/12sin(6t) + 1/96sin(12t), but since anything with sine is going to be 0, it's pretty much 3pi/8 - 0, which is 3pi/8. OMG I kept trying to integrate before switching to the half angle the 2nd time and kept getting pi* some obscure fraction. THANK YOU THANK YOU THANK YOU!

OpenStudy (nikvist):

\[\int\limits_{0}^{\pi}\sin^4{3t}\,\,dt\quad,\quad u=3t\quad,\quad dt=du/3\] \[\frac{1}{3}\int\limits_{0}^{3\pi}\sin^4{u}\,\,du=\frac{6}{3}\int\limits_{0}^{\pi/2}\sin^4{u}\,\,du=2\int\limits_{0}^{\pi/2}\cos^0{u}\cdot\sin^4{u}\,\,du=\]\[=2\cdot\frac{1}{2}B\left(\frac{1}{2},\frac{5}{2}\right)=B\left(\frac{1}{2},\frac{5}{2}\right)=\frac{\Gamma\left(\frac{1}{2}\right)\cdot\Gamma\left(\frac{5}{2}\right)}{\Gamma\left(\frac{1}{2}+\frac{5}{2}\right)}=\frac{\Gamma\left(\frac{1}{2}\right)\cdot\Gamma\left(\frac{5}{2}\right)}{\Gamma\left(3\right)}=\] \[=\frac{\sqrt{\pi}\cdot\frac{3}{4}\sqrt{\pi}}{2!}=\frac{3}{8}\pi\]

OpenStudy (anonymous):

wow awesome LaTeX